OpenStudy (anonymous):

Pens are bought at .95 a dozen and sold in groups of four for $2.25. Find the profit on 15 dozen pens.

myininaya (myininaya):

D=.95 4D=2.25 there are 3 4Ds in 15D after that 3 dozen are left over 3(4D)+3(D)=3(2.25)+3(.95)

OpenStudy (anonymous):

Where did the 3 come from?

OpenStudy (anonymous):

3 x 4= 12 +3= 15

myininaya (myininaya):

right

myininaya (myininaya):

now what is 3(2.25)+3(.95)

OpenStudy (anonymous):

6.75 + 2.85= 9.6

myininaya (myininaya):

yes

OpenStudy (anonymous):

So $9.60 is the profit.

OpenStudy (anonymous):

thanks.